300k A B TANK A CONTANS so ky of FREON -12 at Sara, 20°C TANK B is INITIALLY EMPTY. The VALUE is CRACKED OPEW. The VALVE REMATank A contains 50 kg of freon-12 at 0.5Mpa, at 20 C Tank B is initially empty. The valve is cracked open. The valve remains open. The process ends with both tanks in equilibrium at 0 C saturated vapor. Find the mass in each tank at the end, Q, change is s for tanks, change in sec for surrounding and for the universe, Wcv and Wrev.

Thanks!
Answer #1

PAGE 2 * Solution Considering the figure conditions with schneles from the question zook A 15 Refsigerant freon 12 (8-12) TANfrom ided gl Equation PAGE? So iR-RA is in repour state Enthelpi hi= 196.935 kg/kg Sie 0.6999 Kulkyx - In a osope closed sysPAGE 3 =) 0:3086 X73?_2.015 = M12 x 8314 120.91 * 243 MA= 3312kg MA2 mBz MA imiz 16.88 Kg on the control volume - ConsiderinPAGE 1 DSB = - mp, sit mb 282 -0 + 16-88 X0 696 Goch = +11.7485 15/18 -2 Ethiboy Entropy change of onivesse =DSA+AS + DS un =